# How Do Multilingual Language Models Remember Facts?

### v2

@@ Abstract
Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training. Prior research has localized and identified mechanisms behind knowledge recall; however, it has only focused on English monolingual models. The question of how these mechanisms generalize to non-English languages and multilingual LLMs remains unexplored. In this paper, we address this gap by conducting a comprehensive analysis of three multilingual LLMs. First, we show that previously identified recall mechanisms in English largely apply to multilingual contexts, with nuances based on language and architecture. Next, through patching intermediate representations, we localize the role of language during recall, finding that subject enrichment is language-independent, while object extraction is language-dependent. Additionally, we discover that the last token representation acts as a Function Vector (FV), encoding both the language of the query and the content to be extracted from the subject. Furthermore, in decoder-only LLMs, FVs compose these two pieces of information in two separate stages. These insights reveal unique mechanisms in multilingual LLMs for recalling information, highlighting the need for new methodologies—such as knowledge evaluation, fact editing, and knowledge acquisition—that are specifically tailored for multilingual LLMs.
@@
Constanza Fierro∗†   Negar Foroutan   Desmond Elliott   Anders Søgaard University of Copenhagen    EPFL
**footnotetext: Correspondance: Constanza Fierro <c.fierro@di.ku.dk>.
@@ 1 Introduction
To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric models (Meng et al., 2022). Our results (§3) show that EuroLLM and mT5 exhibit strong causal effects for all subject tokens, with EuroLLM exhibiting this in earlier layers and mT5 across all encoder layers. In XGLM, the last subject token has a stronger effect, but with two early sites, unlike English GPT’s single site. Additionally, in all models, MLPs in later layers are decisive in recovering factual information, a contrast to Meng et al. (2022) findings.
Next, we investigate the second question by analyzing the mechanisms described in Geva et al. (2023); namely, the three-step process, where the relation information flows to the last token, then the subject flows, and finally the attention layers extract the object (§4). We find that in multilingual decoder-only models the information flows similarly to monolingual models, but differently for the encoder-decoder model mT5. In terms of the extraction event, in decoder-only LLMs, both feed-forward and attention sublayers contribute to it, unlike English autoregressive models, where attention modules dominate. In mT5, this mechanism is primarily performed by the cross-attention. Overall, our findings indicate that some of the localization (§3) and mechanisms (§4) of fact retrieval in English LLMs generalize to multilingual LLMs, but with key variations.
Finally, to address the third question, we investigate where language plays a role within the three-step process described by Geva et al. (2023), to characterize how and where factual knowledge cross-lingual transfer may occur. Using activation patching (Zhang and Nanda, 2024; Ghandeharioun et al., 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§5). Our results reveal that the last token acts as a Function Vector (FV) Todd et al. (2024), encoding both the relation and the output language, which is then applied to the subject in the context. Crucially, the fact that the FV formed in one language can be used with an input in another language demonstrates that the subject and relation representations are largely language-independent, while the extraction event is language-specific. Furthermore, in decoder-only LLMs (XGLM, EuroLLM), the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated.
@@ 2 Experimental Setup
We focus on a simple form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$. These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.222 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted. For our analysis, we select 10 typologically diverse languages representing various scripts, families, and word orders: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar). See Table 2 for language characteristics.
@@
We analyze decoder-only and encoder-decoder architectures. [^3] For decoder-only models, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for encoder-decoder, we use all the templates. The decoder-only LLMs are XGLM (Lin et al., 2021),
@@
with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers. The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
Data
We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages. These templates are machine translations of the ParaRel English templates (Elazar et al., 2021). [^4] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix B).
To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023). Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match. If multiple templates yield a match, we randomly select one for the analysis. In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix B.1). Table 1 presents the number of examples for which the correct object is predicted. We exclude languages with too few examples from our analysis, namely ko, ja, he, and fa for XGLM, and ja, he, and fa for EuroLLM.
@@ 3 Causal Tracing
We first analyze which hidden states in the model’s computation are more important than others when recalling a fact. Following Meng et al. (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022). Let $\mathbb{P}(o)$ be the probability of the predicted object token, and $LS(o)$ its logit score. We corrupt the input by adding Gaussian noise to the subject tokens, [^5] We follow Meng et al. (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used. and observe the corrupted probability ~$\tilde{\mathbb{P}}(o)$ of the originally predicted token. Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ~$\tilde{\mathbb{P}}_{\text{restored}}(o)$. We study the indirect effect of such component as $\text{IE}_{\mathbb{P}}=\tilde{\mathbb{P}}_{\text{restored}}(o)-\tilde{\mathbb{P}}(o)$, or if using logits $\text{IE}_{LS}=\tilde{LS}(o)-\tilde{LS}_{\text{restored}}(o)$. Specifically, we restore a state by setting $\tilde{h}_{t}^{l}\leftarrow h_{t}^{l}$, where $h_{t}^{l}$ is the hidden state from the clean run and $\tilde{h}_{t}^{l}$ that of the corrupted run; and similarly, we restore the self-attention layers contribution by setting $\tilde{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$). [^6] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021). In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect. We set $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5, to restore $\sim 12\%$ layers. We repeat the corrupted run with restoration ten times with different noise samples, and report the average.
@@
We compare our results in multilingual LLMs to those of Meng et al. (2022), who analyzed the factual recall of GPT-2 XL in English [^7] They used the CounterFact dataset, which is based on ParaRel. Thus, their English data is the same as our data. and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery. [^8] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration. We, however, interpret the late-site attention as aligning with what Geva et al. (2023) referred to as the extraction event (§4).
@@
Figure 1: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top EuroLLM ($w=7$), bottom mT5 ($w=4$). XGLM in Figure 21.
@@
We present the causal analysis plots for each model and language in Appendix C and present our main observations here. [^9] In line with Zhang and Nanda (2024) we find that analyzing the $\text{IE}_{\mathbb{P}}$ overestimates the causal effect of some tokens over others. For example, in XGLM by looking at the $\text{IE}_{\mathbb{P}}$ we would not find an early site, which is present when studying the $\text{IE}_{LS}$ (see Figure 11 vs Figure 9), or in EuroLLM the last subject token would seem more relevant than the other subject tokens (see Figure 15 vs Figure 13). So we base our main observations using $\text{IE}_{LS}$. Our results indicate that the last subject token does not consistently have a stronger effect on information recovery in early MLPs. In EuroLLM and mT5, all subject tokens exhibit similar effects, while in XGLM, the last subject token has a stronger influence. In EuroLLM, the early site spans half of the initial layers, whereas in mT5, it covers all encoder layers. For XGLM, the early site occurs twice: once in the initial layers and again in the middle layers. Regarding the late site, we observe a strong causal effect in MLP layers before the very last layers in all models. The layers where these sites occur are consistent across languages, with only a slight variation in the late site’s endpoint for EuroLLM.
@@
These findings suggest that some, but not all, conclusions from English LLMs generalize to multilingual LLMs. The MLP early site extends across models, and the MLP late site, which was not observed by Meng et al. (2022), was present in all models studied. This has implications for knowledge localization and fact editing methodologies, which should account for both early and late MLPs in multilingual contexts.
@@ 4 Factual Recall Components
[…]
Figure 1 presents the results with plots per language in Appendix D. Firstly, the results show that in each model the information flows fairly similarly for all the languages, and similar patterns to those found in English GPT are present (Geva et al., 2023). On the one hand, the subject information flows to the last token most critically at the later layers in all models. On the other hand, the relation flows throughout all the layers, with EuroLLM and mT5 having more similar curves. Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in the decoder-only models). [^10] We hypothesize this may occur because the last token encodes which sentinel token is being generated, indicating where to fill in the input. Since we see an almost identical curve when the last token cannot attend to the sentinel token in the decoder (Figure 24).
Figure 2: Extraction rates. Left XGLM, right EuroLLM, bottom mT5.
@@
Following Geva et al. (2023), we measure extraction events at each layer. Let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix; the predicted token is $o=\arg\,\max(Eh^{L})$. An extraction event occurs at layer $l$ if $\arg\,\max(Es^{l})=o$ (similarly for $c^{l}$ and $f^{l}$). The extraction rate is the proportion of examples for which an extraction event occurs at a given layer. Figure 2 shows the extraction rates for each model, with a detailed language breakdown in Appendix E.
Our results demonstrate that extraction events can be detected in multilingual LLMs, though rates vary across languages. A key finding is the prominence of MLP modules in object extraction for multilingual decoder-only models (XGLM and EuroLLM). To rule out the possibility that MLPs simply forward extracted objects from preceding attention layers, we measure how often MLPs perform an extraction without prior attention extraction (Figure 28). We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT-2 results (Geva et al., 2023). In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT-2’s behavior. Additionally, in mT5 the extraction events occur in later layers, with peaks in 19 and 22.
The results show that in multilingual LLMs, information flow from non-subject tokens to the final token is less localized compared to English LLMs, possibly because the model processes relational and language information at different stages. Additionally, the extraction mechanism in multilingual LLMs is more complex, involving both attention and MLP modules. This implies that editing techniques should focus not only on early MLPs enhancing subject representations but also on the later MLPs that extract the object.

### v3

@@ Abstract
Large Language Models (LLMs) store and retrieve vast amounts of factual knowledge acquired during pre-training. Prior research has localized and identified mechanisms behind knowledge recall; however, it has only focused on English monolingual models. The question of how these mechanisms generalize to non-English languages and multilingual LLMs remains unexplored. In this paper, we address this gap by conducting a comprehensive analysis of three multilingual LLMs. First, we show that previously identified recall mechanisms in English largely apply to multilingual contexts, with nuances based on language and architecture. Next, through patching intermediate representations, we localize the role of language during recall, finding that subject enrichment is language-independent, while object extraction is language-dependent. Additionally, we discover that the last token representation acts as a Function Vector (FV), encoding both the language of the query and the content to be extracted from the subject. Furthermore, in decoder-only LLMs, FVs compose these two pieces of information in two separate stages. These insights reveal unique mechanisms in multilingual LLMs for recalling information, highlighting the need for new methodologies—such as knowledge evaluation, fact editing, and knowledge acquisition—that are specifically tailored for multilingual LLMs. [^1] https://github.com/constanzafierro/multilingual_factual_memorization
@@
Constanza Fierro   Negar Foroutan   Desmond Elliott   Anders Søgaard Department of Computer Science, University of Copenhagen EPFL
**footnotetext: Correspondence: Constanza Fierro <c.fierro@di.ku.dk>.
@@ 1 Introduction
Figure 1: Aggregated patching results (§6) for EuroLLM. In propositional factual statements, e.g. “Paris is the capital of”, the last token representation contains the function that solves the task. This function is formed in two stages: first, the relation to extract is encoded (green), and then the language is composed into the function (purple). Finally, the function is applied to the subject, and the predicted object is resolved (blue).
@@
To address the first question, we use causal tracing analysis to assess if early MLP modules processing the final subject token are as decisive in multilingual models as in English-centric ones (Meng et al., 2022). Our results (§4) show that EuroLLM and mT5 exhibit strong causal effects for the last subject token—EuroLLM in earlier layers, and mT5 across all encoder layers. Additionally, in all models, both MLPs and MHSAs in later layers play a decisive role in recovering factual information—unlike in Meng et al. (2022), where only MHSAs were active at the late site.
Next, we investigate the second question by analyzing the three-step process described by Geva et al. (2023): the relation information flows to the final token, followed by subject information, and finally, the attention layers extract the object (§5). We find that in multilingual LLMs, subject information flows similarly to monolingual ones, but non-subject token flow differs, and this analysis alone cannot determine the path of relation information. Regarding the final extraction, both feed-forward and attention sublayers contribute in decoder-only LLMs—unlike in English autoregressive LLMs, where attention modules dominate. In mT5, this mechanism is primarily handled by cross-attention. Overall, our findings indicate that some of the localization (§4) and mechanisms (§5) of fact retrieval in English LLMs generalize to multilingual LLMs, but with key variations.
Finally, to address the third question, we investigate where language plays a role within the three-step process, to characterize how and where factual knowledge cross-lingual transfer may occur. Using activation patching (Zhang and Nanda, 2024) we insert the intermediate representation of the last token from an English forward pass into the forward pass of another language (§6). Our results reveal that the last token acts as a Function Vector (FV) Todd et al. (2024), encoding both the relation and the output language, which is then applied to the subject in the context. Crucially, the fact that the FV formed in one language can be used with an input in another language demonstrates that the subject and relation representations are largely language-independent, while the extraction event is language-specific. Furthermore, in decoder-only LLMs, the FV is constructed in two distinct phases: first, it encodes only the relation, and later, the language is incorporated (Figure 1).
@@
2 Related Work
Factual Knowledge Recall
Petroni et al. (2019) analyzed factual knowledge in pre-trained language models using the LAMA dataset, which pairs cloze-test templates with WikiData triplets. [^3] One might object to treating cloze-test performance as indicative of knowledge, given LLM inconsistencies. However, following Fierro et al. (2024), we adopt this view, as LLMs are often consistent and can sometimes justify responses through world models or training data attribution. This was extended to multilingual settings by translating LAMA (Kassner et al., 2021; Jiang et al., 2020). Later, Elazar et al. (2021) evaluated the consistency of such knowledge by manually curating paraphrases of LAMA to create ParaRel. Further work examined factual consistency cross-lingually (Fierro and Søgaard, 2022; Qi et al., 2023), using machine-translated ParaRel templates and WikiData-based subject/object translations to produce mParaRel.
Interpretability on Factual Knowledge Recall
Early studies by Meng et al. (2022) and Geva et al. (2023) mechanistically analyzed knowledge localization and information flow in GPT models (Brown et al., 2020), using English data. Later, Chughtai et al. (2024) showed that answers emerge from summing independent components in GPT and Pythia models, while Sharma et al. (2024) extended these findings to Mamba (Gu and Dao, 2023), also focusing on English.
Activation Patching
Ghandeharioun et al. (2024) introduced Patchscope, a framework for decoding intermediate representations by patching them into forward passes. Dumas et al. (2024) applied this method to last-token representations in a translation task, showing that models encode language-agnostic concepts—consistent with our findings and those of Foroutan et al. (2022). Wang et al. (2024) similarly found that the last token encodes the relation in factual English queries at a specific stage of computation. We extend these findings to a cross-lingual factual recall setting. While Dumas et al. (2024) concluded that models first resolve the output language, we find that models first resolve the query’s relation, then the language. We thus interpret the last token as a function vector (Todd et al., 2024), which encodes the task-specific function—e.g., translating in the case of Dumas et al. (2024). Additionally, we show that in decoder-only models, the FV composes the output language onto the relation function vector.
@@ 3 Experimental Setup
We focus on a simple form of factual knowledge recall, where LMs are tasked with predicting the correct object $o$ for a given subject $s$ and a relation $r$. These $(s,r,o)$ triplets are obtained from WikiData, and natural language templates are used to describe the relation, with placeholders for the subject and object.444 For example, the relation born-in could use the template “[X] was born in [Y]”, where [X] is the subject and [Y] is the object to be predicted. For our analysis, we select 10 typologically diverse languages representing various scripts, families, and word orders: English (en), Spanish (es), Vietnamese (vi), Turkish (tr), Russian (ru), Ukrainian (uk), Japanese (ja), Korean (ko), Hebrew (he), Persian (fa), and Arabic (ar). See Table 2 for language characteristics.
@@
We analyze decoder-only and encoder-decoder architectures. [^5] For decoder-only models, we only use the templates in mParaRel that have the object placeholder at the end of the sentence, while for encoder-decoder, we use all the templates. The decoder-only LLMs are XGLM (Lin et al., 2021), with 7.5B parameters and 32 layers, and EuroLLM (Martins et al., 2024) with 9B parameters and 42 layers; the encoder-decoder mT5-xl (Xue et al., 2021) has with 3.7B parameters and 24 encoder-decoder layers. The pre-training data varies across models: mT5 is pretrained on 101 languages, covering all the languages in our study; XGLM covers 30 languages, excluding uk, he, and fa; while EuroLLM covers 35 languages, excluding vi, he, and fa.
Data
We use the mParaRel dataset (Fierro and Søgaard, 2022), which includes triplets and templates for 45 languages. [^6] We augment the objects in mParaRel, filter out trivial examples, and when there are enough examples, we use a crosslingual subset (see Appendix A).
Figure 2: Average indirect effect (IE) on logit scores, when the subject input is corrupted and some activations are restored to their uncorrupted values. The $y$ axis indicates the layer where the restoration was performed, and the $x$ axis specifies the token(s) over which we average the IE. The MHSA and MLP are restored in windows of 12% consecutive layers. Top GPT2-XL (only English data), bottom EuroLLM (XGLM and mT5 in Figure 11).
To investigate the process of knowledge recall we only consider examples where the model predicts the correct object completion (Meng et al., 2022; Geva et al., 2023). Since mParaRel provides multiple paraphrased templates for each relation, we greedy-decode for every available template corresponding to a given triplet and check for an exact match.
@@
If multiple templates yield a match, we randomly select one for the analysis. In cases where an article or other filler tokens precede the object, we include these tokens in the input text to ensure that when the example is fed into the model for our analysis, the next predicted token is the first token of the object (Implementation details in Appendix A.1). Table 1 presents the number of examples for which the correct object is predicted. We exclude languages with too few examples from our analysis, namely ko, ja, he, and fa for XGLM, and ja, he, and fa for EuroLLM.
@@ 4 Causal Tracing
We first analyze which hidden states in the model’s computation are more important than others when recalling a fact. Following Meng et al. (2022), we trace the causal effects of hidden states using causal mediation analysis (Pearl, 2022). Let $\mathbb{P}(o)$ be the probability of the predicted object token, and $LS(o)$ its logit score. We corrupt the input by adding Gaussian noise to the subject tokens, [^7] We follow Meng et al. (2022) and add $\epsilon\sim\mathcal{N}(0,(3\sigma)^{2})$, with $\sigma$ being the standard deviation of the subjects tokens embeddings from the data used. We repeat the experiment ten times with different noise samples, and report the average. and observe the corrupted probability ~$\tilde{\mathbb{P}}(o)$ of the originally predicted token. Then, we run inference again on the corrupted input, but this time, we restore a specific hidden state in the model and track the probability ~$\tilde{\mathbb{P}}_{\text{restored}}(o)$. We study the indirect effect of such component as $\text{IE}_{\mathbb{P}}=\tilde{\mathbb{P}}_{\text{restored}}(o)-\tilde{\mathbb{P}}(o)$, or if using logits $\text{IE}_{LS}=\tilde{LS}(o)-\tilde{LS}_{\text{restored}}(o)$. Specifically, we restore a state by setting $\tilde{h}_{t}^{l}\leftarrow h_{t}^{l}$, where $h_{t}^{l}$ is the hidden state from the clean run and $\tilde{h}_{t}^{l}$ that of the corrupted run; and similarly, we restore the self-attention layers contribution by setting $\tilde{s}^{l}\leftarrow s^{l}$ for all the attention modules in a window of size $w$ (analogous for $c$ and $f$). [^8] We use windows because, generally, the contributions of the sub-layers are gradual (Geva et al., 2021). In other words, multiple layers contribute the same behavior to the residual, and their sum produces the observed effect. We restore windows of $12\%$ consecutive layers, so $w$=$4$ for XGLM, $w$=$5$ for EuroLLM, and $w$=$3$ for mT5.
@@
We compare our results in multilingual LLMs to those of Meng et al. (2022), who analyzed the factual recall of GPT-2 XL in English and reached two key conclusions: (1) they identified an “early site”, where the MLPs processing the last subject token in the early and middle layers play a crucial role in recovering from input corruption; and (2), they found a “late site”, where the attention modules processing the last token in the later layers also significantly contribute to prediction recovery. [^9] While they consider the late site unsurprising, as it directly precedes the final prediction, this observation primarily applies to hidden state restoration, not necessarily attention restoration. We, however, interpret the late-site attention as aligning with what Geva et al. (2023) referred to as the extraction event (§5).
@@
Figure 3: Attention knockout between the last token and a given set of tokens. Each layer represents the effect of the knockout on a window of $w$ layers. Top is the average, bottom EuroLLM ($w=7$). XGLM and mT5 in Figure 24.
@@
We present the average causal analysis results in Figure 2 (results per language in Appendix B). [^10] In line with Zhang and Nanda (2024) we find that analyzing the $\text{IE}_{\mathbb{P}}$ overestimates the causal effect of some tokens over others. For example, in EuroLLM the last subject token MHSA seem more relevant than the last token MLPs (see Figure 18 vs Figure 16). So we base our main observations using $\text{IE}_{LS}$. Our results indicate that the early site in MLPs—centered on the last subject token—is also present in multilingual LLMs. We observe this causal effect in both EuroLLM and mT5. In mT5, the first subject token also shows causal relevance, though marginally less so, and the early site spans all encoder layers. Across languages, the patterns are highly consistent. Interestingly, we do not observe an MLP early site in XGLM, either for English or in the MHSAs. As for the late site, we find a strong causal effect in both MLP and MHSA layers just before the final layers in all models. The position of the late site is consistent across languages, with only slight variation in its endpoint for EuroLLM.
@@
These findings suggest that some conclusions from English LLMs generalize to multilingual ones, but not all. The early MLP site is consistent across models, while at the late site, both MLPs and MHSA are important in multilingual LLMs—unlike in English LLMs, where MHSA dominates. This has implications for knowledge localization and fact editing, which should consider both early and late MLPs in multilingual contexts.
@@ 5 Factual Recall Components
We use attention knockout (Geva et al., 2023) to analyze how information propagates to the final token. This method involves intervening in the attention mechanism: for XGLM and EuroLLM, we modify the last token’s self-attention, while for mT5, we intervene in the decoder’s cross-attention for the final token. We knock out attention connections by setting the attention scores to zero between the final token and a token set $\{t\}$, which can be: subject tokens, non-subject tokens, or the last token itself. Following Geva et al. (2023) we apply this intervention across consecutive layers within a window of size $w$ centered on layer $l$, covering $18\%$ of the total layers ($w$=$6$ for XGLM, $w$=$7$ for EuroLLM, and $w$=$4$ for mT5). The relative probability change is given by (~$(\tilde{\mathbb{P}}(o)-\mathbb{P}(o))/\mathbb{P}(o)$, where ~$\tilde{\mathbb{P}}(o)$ is the probability of the originally predicted token $o$ after attention knockout .A significant drop indicates that the knocked-out tokens contribute critically to the final token’s prediction at that layer.
In Figure 3 we present average results, plots per language can be found in Appendix C. The results show that in each model the information flows fairly similarly for all the languages, and somewhat similar patterns to those in English GPT2-xl. On the one hand, the subject information flows to the last token most critically at the later layers in all models. On the other hand, the non-subject tokens information flows throughout all the layers, with EuroLLM and mT5 having more similar curves. Nevertheless, we note two differences from the conclusions reached with English GPT (Geva et al., 2023): (1) the propagation of non-subject tokens to the last token does not strictly precedes the subject propagation, in XGLM the drop is the highest in the earlier layers but it continues to be important until the end, and in mT5 and EuroLLM this information has two peaks, in the early and later layers; and (2) the last token of mT5 encodes critical information that flows from one layer to the next through the cross-attention (as opposed to the negligible flow of the last token in decoder-only models). [^11] We hypothesize this may occur because the last token encodes which sentinel token is being generated, indicating where to fill in the input. Since we see an almost identical curve when the last token cannot attend to the sentinel token in the decoder (Figure 27).
Figure 4: Extraction rates. Fraction of examples where the top-1 token under the vocabulary projection matches the final output token.
@@
Following Geva et al. (2023), we measure extraction events at each layer. Let $h^{l}$ be the representation of the last token at layer $l$, and let $E$ be the embedding matrix; the predicted token is $o=\arg\,\max(Eh^{L})$. An extraction event occurs at layer $l$ if $\arg\,\max(Es^{l})=o$ (similarly for $c^{l}$ and $f^{l}$). The extraction rate is the proportion of examples for which an extraction event occurs at a given layer.
Our results demonstrate that extraction events can be detected in multilingual LLMs (Figure 4), though rates vary across languages (Appendix D). [^12] This variance may result from the strict criteria applied during the extraction event analyses, where only top-1 matches were considered. A key finding is the prominence of MLP modules in object extraction for multilingual decoder-only models (XGLM and EuroLLM). To rule out the possibility that MLPs simply forward extracted objects from preceding attention layers, we measure how often MLPs perform an extraction without prior attention extraction (Figure 31). We find that MLPs indeed perform the extraction for most languages, though in English, attention modules dominate, aligning with GPT2-xl results. In mT5, cross-attention layers drive the extraction, with MLPs playing a secondary role, resembling GPT’s behavior. Additionally, in mT5 the extraction events occur in later layers, with peaks in 19 and 22.
The results show that in multilingual LLMs the extraction mechanism is more complex, involving both attention and MLP modules. This implies that editing techniques should focus not only on early MLPs enhancing subject representations but also on the later MLPs that extract the object.